OpenStudy (anonymous):

if I have this equation: -4jA + (6+4j)B - 60 + 50<30 = 0 is it the same as this? -4jA + (6+4j)B = -60 + 50<-150

OpenStudy (anonymous):

@optiquest

OpenStudy (anonymous):

I dont think, you should move them around like regular

OpenStudy (anonymous):

let's say that equation is from a mesh loop. I have to transfer the voltages at the other side of the equation. i think it should be -4jA + (6+4j)B = 60+50<-150 is it?

OpenStudy (anonymous):

-4jA + (6+4j)B =60 - 50<30 imo

OpenStudy (anonymous):

60-50<30 is math error in my calculator XD

OpenStudy (anonymous):

you try with parenthesis? works in mine

OpenStudy (anonymous):

it is but my book does not agree. XD

OpenStudy (anonymous):

i get 30.06<-56.26

OpenStudy (unklerhaukus):

-4jA + (6+4j)B - 60 + 50 < 30 adding 60 to both sides (of the in-equation) -4jA + (6+4j)B + 50 < 30 +60 minus-ing 50 -4jA + (6+4j)B < 30 + 60 - 50 simplifying -4jA + (6+4j)B < 40

OpenStudy (anonymous):

its angle not less than

OpenStudy (unklerhaukus):

oh, right